# Uncertainty-Aware Collaborative System of Large and Small Models for Multimodal Sentiment Analysis

### v1

@@ Abstract
The advent of Multimodal Large Language Models (MLLMs) has significantly advanced the state-of-the-art in multimodal machine learning, yet their substantial computational demands present a critical barrier to real-world deployment. Conversely, smaller, specialized models offer high efficiency but often at the cost of performance. To reconcile this performance-efficiency trade-off, we propose a novel Uncertainty-Aware Collaborative System (U-ACS) that synergistically orchestrates a powerful MLLM (e.g., HumanOmni) and a lightweight baseline model for multimodal sentiment analysis. The core of our system is an uncertainty-driven cascade mechanism, where the efficient small model first acts as a rapid filter for all input samples. Only those samples yielding high predictive uncertainty, thereby indicating greater difficulty, are selectively escalated to the MLLM for more sophisticated analysis. Furthermore, our system introduces advanced strategies to handle ambiguous or conflicting predictions, including weighted averaging for predictions of similar polarity and a prompt-based cross-verification to resolve conflicting predictions when both models exhibit high uncertainty. This sample-difficulty-aware approach allows for a dynamic allocation of computational resources, drastically reducing inference costs while retaining the high accuracy of MLLM. Extensive experiments on benchmark datasets demonstrate that our proposed method achieves state-of-the-art performance, while requiring only a fraction of the computational resources compared to using a standalone MLLM.
Introduction
@@
The landscape of multimodal learning (Baltrušaitis, Ahuja, and Morency 2019) has been profoundly reshaped by the emergence of powerful Multimodal Large Language Models (MLLMs) such as HumanOmni (Zhao et al. 2025). Multimodal Sentiment Analysis (MSA), leveraging acoustic, textual, and visual information to predict human sentiment intensities and emotional states (Poria et al. 2017), can also benefit from the capability of MLLMs. With their vast parameter counts and sophisticated reasoning capabilities, MLLMs demonstrates an unparalleled ability to interpret complex, nuanced human expressions across textual, visual, and acoustic modalities. However, this remarkable accuracy comes at a significant cost. The substantial computational overhead associated with MLLMs poses a critical barrier to their deployment in many practical, real-world scenarios. This overhead manifests as high inference latency, significant memory requirements, and considerable energy consumption, particularly problematic where resources are constrained or real-time responses are paramount.
At the other end of the spectrum, there are smaller and specialized models (Wang et al. 2025a; Yu et al. 2021). These lightweight architectures are characterized by their computational efficiency, offering rapid inference speeds and minimal resource footprints. Yet, this efficiency is often achieved by sacrificing the deep, generalizable understanding inherent to MLLMs, resulting in a performance ceiling, particularly when confronted with ambiguous or out-of-distribution samples. This dichotomy presents a challenging trade-off for the field: a choice between the high accuracy of costly large models and the high efficiency of less capable small models. The pursuit of a solution that marries the strengths of both paradigms remains a significant and open research question.
Figure 1: Advantages and limitations of existing approaches: small models offer efficiency but lack accuracy, while MLLMs provide high accuracy at heavy computational cost.
To bridge this gap, we introduce a novel Uncertainty-Aware Collaborative System (U-ACS) that intelligently orchestrates the operations of a powerful MLLM and an efficient small baseline model. As conceptually illustrated in Figure 1, our approach is not a static ensemble but a fluid, sample-aware cascade designed to achieve optimal performance with minimal average computational cost. The core principle is to allocate computational resources judiciously, reserving the analytical power of the MLLM exclusively for samples that genuinely require it.
The cornerstone of the proposed system is an uncertainty-guided routing mechanism. We posit that a model’s predictive uncertainty is a reliable proxy for sample difficulty (Tellamekala et al. 2024; Gao et al. 2024). Consequently, both the small models and MLLMs are trained to co-produce not only a sentiment prediction but also a quantifiable validation-calibrated measure of their own uncertainty. The inference process begins with the efficient small model, which acts as a rapid initial filter for all incoming samples. For a given sample, if the small model exhibits low uncertainty (i.e., high confidence), its prediction is accepted, and the process terminates. This ‘fast path’ effectively handles the majority of straightforward cases with minimal computational expense.
Only when the small model’s uncertainty exceeds a predefined threshold, indicating a ‘hard’ sample, is the inference process escalated to the MLLM. This selective escalation ensures that our most potent computational asset is used purposefully. Upon processing the sample, the MLLM also provides its prediction and uncertainty. If the MLLM is confident in its assessment, its superior judgment is adopted as the final result. However, the most intricate scenarios arise when the MLLM is also uncertain where both models produce predictions that require further deliberation. To resolve these problem, we introduce a sophisticated collaboration logic. When both models are confident and their predictions are similar in magnitude (e.g., both positive or both negative with comparable values), we employ a weighted average based on their uncertainty scores to refine the final prediction. In cases where the models produce conflicting sentiment polarities (e.g., one positive, one negative), we utilize a prompt-based cross-verification strategy, where both model’s predictions and uncertainties are incorporated into the MLLM’s prompt to stimulate re-evaluation and collaborative reasoning.
This hierarchical and adaptive approach allows our system to dynamically navigate the performance-cost landscape, delivering a solution that is both highly accurate and computationally tractable. Our main contributions can be summarized as follows:
@@
  • We propose a novel U-ACS that synthesizes a large, powerful MLLM with a small, efficient model to resolve the performance-versus-efficiency dilemma in MSA.
@@
  • We introduce a sophisticated collaboration logic, including prompt-based cross-verification, to resolve predictive conflicts and ambiguities between two models.
  • Through extensive experiments, we demonstrate that our system achieves state-of-the-art performance on multiple MSA datasets while operating at a fraction of the computational cost of a standalone MLLM approach.
Related Work
Multimodal Sentiment Analysis
Early MSA research focused on feature fusion and representation learning techniques, ranging from simple concatenation and tensor-based methods (Zadeh et al. 2017; Liang et al. 2019) to attention-based mechanisms for capturing inter-modal dependencies (Tsai et al. 2019; Mai, Xing, and Hu 2021). In parallel, recent approaches have explored graph neural networks for modeling speaker-utterance relationships in conversational settings (Tu et al. 2024; Wang et al. 2025b) and self-supervised frameworks for handling incomplete or noisy data (Li et al. 2024). To improve efficiency beyond traditional Transformers, State Space Models like MSAmba (He et al. 2025) have demonstrated effective long-range modeling with lower computational complexity. However, these methods typically operate as monolithic systems, either prioritizing maximum performance or efficiency, without dynamic resource allocation mechanisms.
@@
Large Language Models (LLMs) have catalyzed significant advances in MSA (Yang et al. 2024). These methods often use lightweight adapters like MSE-Adapter (Yang, Dong, and Qiang 2025) to enable multimodal capabilities in frozen LLMs while reducing training costs. More recently, Multimodal Large Language Models (MLLMs) have emerged, which can directly handle multimodal signals and project them into the feature space of LLMs and achieve better results (Zhao et al. 2025; Wu et al. 2025). While effective, these approaches invoke MLLMs for every sample, remaining computationally expensive for practical deployment.
Collaboration of Large and Small Models
The concept of combining large, high-capacity models with small, efficient ones to optimize the performance-cost trade-off is an emerging trend in the broader machine learning community (Min et al. 2024; Chen, Zhao, and Han 2025). These model cascades use small models for easy cases and large models for complex inputs, successfully applied in computer vision and NLP tasks requiring low latency.
In the MSA domain, this collaborative paradigm is still nascent. The most related concepts can be found in knowledge distillation frameworks, where a larger ”teacher” model’s knowledge is transferred to a smaller ”student” model during training (Li et al. 2024). However, these methods focus on creating a single, more compact model for inference, rather than using both models dynamically during inference. To the best of our knowledge, a dynamic, uncertainty-driven cascade that synthesizes a massive MLLM with a specialized small model for MSA has not been explored. Our work is the first to propose such a framework, explicitly designed to balance the unparalleled accuracy of models like HumanOmni with the practical efficiency required for real-world applications.
Methodology
Problem Definition
@@
The goal of MSA is to predict the sentiment intensity from given multimodal inputs. In the proposed U-ACS, we simultaneously incorporate three modalities: Text (T), Vision (V), and Audio (A). Each modality is represented by a tensor $\hat{X}^{m}\in\mathbb{R}^{N_{m}\times d_{m}}$, where $N_{m}$ denotes the sequence length, $d_{m}$ is the embedding dimension, and $m\in\{T,V,A\}$ indicates the specific modality type.
Figure 2: Three stages compose our U-ACS. The efficient baseline model acts as a primary filter, escalating only high-uncertainty samples to the MLLM (HumanOmni), followed by collaborative processing to resolve prediction conflicts.
Overall Architecture
As shown in Figure 2, our proposed U-ACS synergistically integrates a lightweight small model (Uncertainty-aware Baseline Model, UBM) with a powerful MLLM through an uncertainty-guided cascade, consisting of three stages: i) In the first stage, the small model quickly processes all input samples, efficiently filtering out simple cases by making high-confidence predictions with minimal computational cost. ii) In the second stage, samples that exhibit higher predictive uncertainty are escalated to the LoRA-adapted MLLM, which performs more in-depth, modality-specific analysis to improve prediction accuracy. iii) In the third and final stage, the outputs from both models are combined through a collaborative reasoning mechanism, leveraging their complementary strengths to generate robust and reliable predictions, especially for the most complex cases.
Small Model Filtering
The small model acts as the primary inference engine of the system, optimized for efficient multimodal learning and high-throughput processing. It delivers fast sentiment predictions for the majority of samples while generating reliable uncertainty estimates to guide sample routing within the cascade.
Unimodal Learning
The model processes each modality using dedicated encoders. For the text modality, we extract the representation at the first time step:
$\begin{split}\bm{\hat{X}}_{t}&=TextEncoder(\bm{U}_{t};\theta_{t})\in\mathbb{R}^{T\times d_{t}},\\ \bm{x}_{t}&=\bm{\hat{X}}_{t}[0,:]\in\mathbb{R}^{d_{t}}\end{split}$ (1)
For the acoustic and visual modalities, the last time step representation is extracted:
$\begin{split}\bm{\hat{X}}_{a}&=AudioEncoder(\bm{U}_{a};\theta_{a})\in\mathbb{R}^{T\times d_{a}},\\ \bm{\hat{X}}_{v}&=VisualEncoder(\bm{U}_{v};\theta_{v})\in\mathbb{R}^{T\times d_{v}},\\ \bm{x}_{a}&=\bm{\hat{X}}_{a}[-1,:],\quad\bm{x}_{v}=\bm{\hat{X}}_{v}[-1,:]\end{split}$ (2)
where $\bm{x}_{t}$, $\bm{x}_{a}$, and $\bm{x}_{v}$ represent the extracted feature vectors for the text, audio, and visual modalities, respectively.
Multimodal Fusion and Prediction
The feature vectors $\bm{x}_{t}$, $\bm{x}_{a}$, and $\bm{x}_{v}$ are concatenated and fused for prediction:
$\begin{split}\bm{h}_{m}&=Fusion(\bm{x}_{t}\|\bm{x}_{a}\|\bm{x}_{v})\in\mathbb{R}^{d_{m}},\\ \hat{y}_{s}&=Predictor(\bm{h}_{m})\end{split}$ (3)

### v2

@@ Abstract
Multimodal Large Language Models (MLLMs) have notably enhanced the performance of Multimodal Sentiment Analysis (MSA), yet their massive parameter scale leads to excessive resource consumption in training and inference, severely limiting model efficiency. To balance performance and efficiency for MSA, this paper innovatively proposes a novel Uncertainty-Aware Collaborative System (U-ACS) that integrates Uncertainty-aware Baseline Model (UBM) with MLLMs. U-ACS operates in three stages: First, all samples are processed by the UBM, retain high-confidence samples and forward low-confidence samples to the MLLM. Notably, to address the challenge that continuous outputs of regression tasks hinder uncertainty calculation, we innovatively convert the continuous sentiment label prediction task to a classification task, enabling a more accurate calculation of entropy and uncertainty. Second, the MLLM performs initial process. In this stage, high-confidence samples or low-confidence samples whose predictive sentiment polarity matches that of the UBM are deemed acceptable, while unqualified samples are forwarded for further processing. Finally, the MLLM performs secondary inference on remaining low-confidence samples using prompts augmented with prior rounds predictions as references. By aggregating results from the three stages, U-ACS preserves high MSA prediction accuracy while drastically boosting efficiency via offloading most simple samples to the UBM and minimizing MLLM processing volume. Extensive experiments verify that U-ACS maintains superior performance while significantly reducing computational overhead and resource consumption.
Index Terms:
Multimodal Large Language Model, Multimodal Sentiment Analysis, Collaborative of Large and Small Models, Uncertainty
I Introduction
In the real world, there exists a vast amount of multi-source and heterogeneous multimodal data. Such information carriers, which integrate different modalities including text, images, audio, and video, serve as the core medium for humans to perceive and cognize the entire objective world. Consequently, Multimodal Sentiment Analysis (MSA), which enables the comprehensive analysis of users’ emotional inclinations and opinion trends by leveraging the complementary information in multimodal data, has increasingly become a focal point of academic and industrial research [34, 47, 29].
@@ I Introduction
The landscape of multimodal learning [1] has been profoundly reshaped by the emergence of powerful Multimodal Large Language Models (MLLMs) such as HumanOmni [64] and VideoLLaMA2 [5]. Leveraging acoustic, textual, and visual information to predict human sentiment intensities and emotional states [34, 22], MLLMs deliver enhanced performance in MSA. With their vast parameter counts and sophisticated reasoning capabilities, MLLMs demonstrates an unparalleled ability to interpret complex, nuanced human expressions. However, this remarkable accuracy comes at a significant cost that manifests as high inference latency, significant memory requirements, and considerable energy consumption, particularly problematic where resources are constrained or real-time responses are paramount.
In addition to MLLMs, academia and industry have also proposed a series of small-parameter specialized models [44, 59] to address MSA tasks. Such lightweight models boast prominent advantages of high computational efficiency, fast inference speed, and low resource footprint, making them well-suited for deployment scenarios with limited computing power such as edge devices. However, their high efficiency is often achieved at the expense of overall inference performance: when confronted with samples featuring imbalanced data distribution, complex modality information, or missing data in partial modalities, the sentiment recognition accuracy and robustness of lightweight models will experience a significant decline, and their processing capability will be severely constrained. Based on the inherent strengths and weaknesses of these two types of models, a key challenge has emerged in the MSA field: how to significantly improve the operational efficiency of models while maintaining their original inferential performance for MSA tasks? Therefore, exploring technical solutions that can integrate the advantages of both paradigms remains an important and open research issue in this domain.
Figure 1: Advantages and disadvantages of small models and MLLMs. Small models exhibit superior performance on efficiency, whereas large models demonstrate more advantages in performance.
Figure 1 systematically summarizes the core strengths and weaknesses of lightweight small models and MLLMs. It can be observed that there is a significant gap between the two types of models in terms of computational efficiency and inference speed: small models possess distinct lightweight advantages, whereas MLLMs exhibit high resource consumption. In terms of task performance, although MLLMs outperform small models overall, the sentiment analysis accuracy of small models still falls within an acceptable range. In view of the characteristics and shortcomings of these two types of models, this paper proposes a novel Uncertainty-Aware Collaborative System (U-ACS). This system innovatively achieves hierarchical collaboration between Uncertainty-aware Baseline Model (UBM) and MLLM, and realizes dynamic and adaptive allocation of computing resources between the two types of models through the design of a three-stage progressive processing workflow.
Specifically, U-ACS quantifies uncertainty [40, 11] to evaluate the model’s predictive performance on specific samples. In the first stage of U-ACS, all samples are fed into the UBM for processing, with uncertainty scores computed for each sample. Samples with low uncertainty have high-confidence predictions from the UBM preserved; those with high uncertainty are forwarded to the MLLM for follow-up analysis. Notably, as regression tasks produce continuous outputs that hinder the calculation of corresponding uncertainty, we innovatively convert continuous sentiment label prediction task to a classification task, and build a three-class classifier to predict the discrete sentiment label (positive, negative and neutral). This operation allows us to accurately calculate sample entropy, which is then used to measure sample uncertainty.
Another two stages are operated within the MLLM module. In the second stage, all samples with uncertainty exceeding the threshold in the UBM stage are fed into the MLLM for the first round of verification. Samples are categorized into three types according to their uncertainty scores and the consistency of sentiment polarity: high-confidence samples with low uncertainty, low uncertainty samples whose predicted sentiment polarities are consistent between UBM and MLLM, and high uncertainty samples with inconsistent sentiment polarities predicted by the two models. Among these, the prediction results of the first two types of samples can be directly adopted. The first type has high reliability due to sufficient confidence, while the second type, despite insufficient confidence, has consistent polarity judgments from the small and large models, making it unlikely that subsequent predictions will alter the core conclusion. Thus, only the third type of samples proceeds to the second round of in-depth inference in the MLLM. In the third stage, we embed the sentiment intensity predictions of these samples from the small model stage and the first-round MLLM results into the prompt as multi-dimensional reference bases, enabling the MLLM to make more accurate predictions for the remaining samples. Finally, by integrating the high-confidence results from the small model stage, the acceptable results from the first MLLM round, and the final discrimination results from the second MLLM round, we can obtain complete multimodal sentiment analysis results for all samples.
This hierarchical processing mechanism can significantly improve the operational efficiency of the model while ensuring the overall performance of MSA tasks, thereby effectively balancing the inherent trade-off between performance and efficiency. Our main contributions can be summarized as follows:
@@
  • We propose a novel U-ACS that synthesizes MLLM with UBM to resolve the performance-versus-efficiency dilemma in MSA. To the best of our knowledge, this is the first small-large model collaboration scheme proposed for the MSA task.
@@
  • Through extensive experiments, we demonstrate that our strategy achieves state-of-the-art performance on part of the key benchmarks across multiple MSA datasets, while also significantly improving the overall inference efficiency and reducing computational overhead substantially compared to MLLM.
II Related Work
II-A Multimodal Sentiment Analysis
The core task of MSA [10] is to fuse information from multimodalities such as text, images, audio, and video, so as to automatically identify and determine the sentiment tendency, intensity, or category implied in the target object (i.e., sentences, paragraphs, or video clips). Existing studies have predominantly focused some research directions in multimodal learning: developing sophisticated fusion mechanisms to characterize cross-modal interactive relationships [34, 35, 18, 41, 27, 28]; designing robust algorithmic frameworks to mitigate the adverse impacts of noisy modalities [20, 46, 19, 25, 26]; constructing weakly supervised and self-supervised learning paradigms to alleviate the constraint of scarce training data [31]; and devising effective optimization algorithms to bridge inter-modality gaps while regularizing the learning process of unimodal data distributions [13, 39, 49]. However, all the aforementioned core research endeavors are confronted with a critical challenge: the high complexity of MSA tasks necessitates the adoption of MLLMs to ensure satisfactory performance. Meanwhile, such task complexity also drives the design of sophisticated algorithms, which in turn leads to inefficient inference of MLLMs and substantial computational resource consumption. To improve efficiency beyond traditional Transformers, State Space Models like MSAmba [14] have demonstrated effective long-range modeling with lower computational complexity. However, these methods typically operate as monolithic systems, either prioritizing maximum performance or efficiency, without dynamic resource allocation mechanisms.
@@ II-A Multimodal Sentiment Analysis
Meanwhile, Large Language Models (LLMs) have catalyzed significant advances in MSA [54, 30]. These methods often use lightweight adapters like MSE-Adapter [56] to enable multimodal capabilities in frozen LLMs while reducing training costs. More recently, MLLMs have emerged, which can directly handle multimodal signals and project them into the feature space of LLMs and achieve better results [64, 48]. While effective, these approaches invoke MLLMs for every sample, remaining computationally expensive for practical deployment.
II-B Collaboration of Large and Small Models
An emerging trend across the machine learning field is the integration of large, high-capacity models and small, efficient models, which aims to optimize the trade-off between model performance and computational cost [32, 4]. Such cascaded systems are designed to route straightforward inputs to lightweight models, with large models only invoked for handling challenging cases; this design has been validated effectively in low-latency computer vision and Natural Language Processing (NLP) tasks [2].
In the field of MSA, research on the collaboration mechanism between large and small models is still in the preliminary exploration stage, where existing methods mostly adopt a single large model or small model to complete tasks independently. In contrast, in other relevant machine learning fields, to facilitate the widespread deployment of models on resource-constrained devices, the collaboration mechanism between large and small models has gradually become a research hotspot jointly concerned by academia and industry. Chen et al. [4] categorized the collaboration patterns of large and small models into five major types: Pipeline Collaboration [24, 63, 21, 3, 51], Auxiliary/Enhancement Collaboration [37, 17, 50, 6], Knowledge Distillation-Driven Collaboration [52, 51, 15, 12], and Integration/Fusion Collaboration [33, 24, 38, 8, 43], Hybrid/Routing Collaboration [24, 45, 42]. Models characterized by distinct scales and performance capabilities are deployed on a task-dependent basis. A binary classifier is incorporated into HybridLLM [57] to assess query difficulty, with the system then directing queries to different models in accordance with the generated predictions. The Collaborative Inference with Token-lEvel Routing (CITER) framework [65], a typical paradigm of routing collaboration, is adopted by this framework, with non-critical tokens being dispatched to the Small Language Model (SLM) for improved efficiency and critical tokens being transmitted to the LLM to ensure generation quality. Such fine-grained routing enables dynamic tuning of the efficiency-quality balance within the system. Consistent with the above paradigm, our strategy is likewise classified as a Hybrid/Routing Collaboration approach.
Different from previous studies, our strategy prioritizes the improvement of model efficiency rather than the pursuit of pure performance gains. We aim to explore a viable paradigm for balancing efficiency and performance, and thus focus on developing an approach that can enhance efficiency while preserving the original performance level. To achieve this goal, this paper introduces a dedicated strategy that can fulfill this objective in MSA scenarios.
III Model Architecture
In this section, we elaborate on the details of our proposed U-ACS framework, whose overall architecture is illustrated in Fig. 2. The detailed algorithmic implementation of each component is presented in the subsequent subsections.
III-A Task Definition
@@
The goal of MSA is to predict the sentiment intensity from given multimodal inputs. In this task, we consider three different modalities, i.e., text ($t$), visual ($v$), and acoustic ($a$). The three modality is represented by sequences $\bm{U}_{t}\in\mathbb{R}^{N_{t}\times d_{t}}$, $\bm{U}_{v}\in\mathbb{R}^{N_{v}\times d_{v}}$, $\bm{U}_{a}\in\mathbb{R}^{N_{a}\times d_{a}}$,where $N_{m}$ denotes the sequence length, $d_{m}$ is the embedding dimension, and $m\in\{t,v,a\}$ indicates the specific modality type. The input feature sequence $\bm{U}_{m}\in\mathbb{R}^{N_{m}\times d_{m}}$ is mapping to unimodel representation $\bm{x}_{m}\in\mathbb{R}^{d_{m}}$ through the unimodel learning network $F^{m}$. The full structural details of unimodel learning network $F^{m}$ and fusion network $F^{f}$ in the subsequent subsections.
Figure 2: The pipeline of the U-ACS. The diagram depicts the three-stage pipeline of U-ACS, where the first stage relies on the UBM, and the latter two stages are centered on the MLLM. The workflow is structured to leverage the complementary advantages of UBM and MLLM for enhanced MSA efficiency.
III-B Overall Architecture
U-ACS establishes a synergistic paradigm that integrates UBM with MLLMs. Leveraging the complementary strengths of the two model types, where UBM excel in computational efficiency and MLLMs dominate in sentiment analysis performance, we design a tiered processing pipeline: UBM are tasked with predicting the sentiment intensity of easy samples, and difficult samples are forwarded to MLLMs for in-depth inference. This division of labor ensures that the framework maintains high sentiment analysis performance while significantly boosting overall inference efficiency. We will introduce the pipeline in UBM and MLLM separately.
III-C Pipeline of UBM
This subsection presents a comprehensive overview of the UBM, which is structured as follows: first, we decompose its core architecture, elaborate on the module functions and connection mechanisms of the encoder, fusion network, and predictor in sequence; second, we detail the uncertainty quantification method for the model’s outputs, which provides a theoretical basis for subsequent sample selection in the small-large model collaborative framework; finally, we present the mathematical formulation of the loss function adopted in the model training phase and offer corresponding interpretations.
1) Model-specific Encoders
Language Encoder: For textual input processing, we adopt the BERT-base model [7] as the backbone encoder, where the English variant (bert-base-uncased) and Chinese variant (bert-base-chinese) are dynamically selected based on the language of the target dataset. Text sequences are first tokenized using the official BERT tokenizer, and the tokenized outputs are fed into the pre-trained BERT model to generate contextualized text representations:
$\begin{split}\bm{H}_{t}&=\text{BERT}(\bm{U}_{t})\in\mathbb{R}^{N_{t}\times d_{bert}},\\ &\bm{x}_{t}=\bm{H}_{t}[0,:]\in\mathbb{R}^{d_{bert}}\end{split}$ (1)
where $\bm{U}_{t}$ denotes the tokenized input of a single text sample, $\bm{H}_{t}$ represents the contextualized hidden state sequence output by BERT for the single text sample, with its dimension $\mathbb{R}^{N_{t}\times d_{\text{bert}}}$, $\bm{x}_{t}$ is the final text representation extracted from the 0-th token (i.e., the [CLS] token) of $\bm{H}_{t}$, which aggregates the global semantic information of the entire text sequence and has a dimension of $\mathbb{R}^{d_{\text{bert}}}$. Notably, the dimension $d_{\text{bert}}=768$ is kept consistent across all datasets in our experiments for single-sample text representation.
Audio Encoder: Following Zadeh et al. [60] and Yu et al. [58], we ues pre-train ToolKits to extract the initial features in audio and visual modalities. For acoustic feature processing, following Self-MM [59], we utilize a lightweight LSTM-based subnet to capture temporal dependencies, which has few parameters:
$\begin{split}\bm{H}_{a}=&\text{LSTM}(\bm{U}_{a};\theta_{a})\in\mathbb{R}^{N_{a}\times d_{a}},\\ &\bm{x}_{a}=\bm{H}_{a}[-1,:]\in\mathbb{R}^{d_{a}}\end{split}$ (2)
where $\bm{U}_{a}$ represents the input audio features, $\theta_{a}$ represents the full set of learnable parameters of the LSTM subnet, and $\bm{H}_{a}\in\mathbb{R}^{N_{a}\times d_{a}}$ is the end-state hidden vector output by the LSTM, which aggregates the global temporal dependencies of the entire audio sequence. $\bm{x}_{a}\in\mathbb{R}^{d_{a}}$ is the final audio representation with $d_{a}=16$.
$\begin{split}\bm{H}_{v}=&\text{LSTM}(\bm{U}_{v};\theta_{v})\in\mathbb{R}^{N_{v}\times d_{v}},\\ &\bm{x}_{v}=\bm{H}_{v}[-1,:]\in\mathbb{R}^{d_{v}}\end{split}$ (3)
where $\bm{U}_{v}$ represents the input audio features, $\theta_{v}$ represents the full set of learnable parameters of the LSTM subnet, and $\bm{H}_{v}\in\mathbb{R}^{N_{v}\times d_{v}}$ is the end-state hidden vector output by the LSTM, which aggregates the global temporal dependencies of the entire audio sequence. $\bm{x}_{v}\in\mathbb{R}^{d_{v}}$ is the final audio representation with $d_{v}=32$.
2) Multimodal Feature Fusion
The core function of the fusion network is to perform modal concatenation on the three types of unimodal representation, merging text, audio, and visual features into a holistic fused representation:
$\begin{split}&\bm{x}_{concat}=\bm{x}_{t}\|\bm{x}_{a}\|\bm{x}_{v},\\ \bm{h}_{m}&=Fusion(\bm{x}_{concat})\in\mathbb{R}^{d_{m}}\end{split}$ (4)